OpenStudy (anonymous):

when i use elimination to solve a system of equations, what do i do if there is nothing to eliminate? ex. 5x+4y=-6 3x-3y=4

OpenStudy (anonymous):

you can always eliminate ex multiply the first line by 3 and the second by -5

OpenStudy (anonymous):

oh ok, i didn't know that you could multiply by two different numbers

hero (hero):

Alternative Hint: \(y = -\large\frac{6 + 5x}{4}\) \(y = -\large\frac{3x - 4}{3}\)

hero (hero):

Solve each for y, then set y = y and cross multiply

hero (hero):

So \(\frac{6 + 5x}{4} = \frac{4 - 3x}{3}\) 3(6 + 5x) = 4(4-3x) then continue solving for x